# Surface Plasmon Polaritons: Creation Dynamics and Interference of Slow and Fast Propagating SPPs at a Temporal Boundary

### v1

@@ I Introduction
Lately, there has been a resurgence in this research topic [7, 6, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 10, 24, 25, 26, 27, 11, 28, 29, 12, 30, 31, 32, 33, 34, 35, 36, 37, 38], driven by the potential for the applications in plasmonics (nano-photonics, nano-optics, photonic metamaterials), such as magnet-free nonreciprocity [15, 29], temporal aiming [21, 39], and extreme energy transformations [28], to name a few. There has been some work on time-varying media systems that support SPPs [16, 18, 22]; however, these deal with existing propagating SPPs encountering a temporal boundary. To the best of our knowledge, there has not been work on dipole excitation of SPPs in a time-varying system. This is the topic of this work, where such a system allows for the analysis of the interaction among dipole excitations, where we have the ability to consider the interactions at the moment of SPP creation (e.g., as a time-varying system changes from one that does not support SPPs to one that supports SPP propagation).
@@
In the subsequent sections we establish the theoretical framework for the specific configuration of a time-varying media system that supports SPPs, in at least one temporal region, or both, excited by a dipole source, where the media configuration suddenly changes at a temporal boundary. We first state Maxwell’s equations in the Laplace transform domain, derive the inhomogenous electric field wave equation in the Laplace transform domain for generic media, i.e., linear, inhomogeneous, anisotropic, temporally dispersive (temporally nonlocal), and spatially nondispersive (local) media, and then illustrate the use of Green’s functions to solve the inhomogenous wave equation to obtain the electric field.
@@
The model for this configuration, seen in Fig. 1, defines two temporal regions (one temporal boundary).
Figure 1: Model for the general case of a single interface between two different materials ($\mu_{2}$,$\varepsilon_{2}$ for $z>0$, $\mu_{1}$,$\varepsilon_{1}$ for $z<0$) with an electric dipole point source, $\widehat{\mathbf{z}}j_{zs}$, located along the interface surface, where we consider two temporal regions (one temporal boundary). The double subscript notation is described in the text. Additionally, $q_{x}$, $q_{y}$, and $\phi$ define the coordinate system in the momentum space.
@@ II-A Maxwell’s Equations in the Laplace Transform Domain
×$\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-s\mu_{0}\underline{\boldsymbol{\mu}}_{r}(\mathbf{r},s)\cdot\mathbf{H}(\mathbf{r},s)$ (1)
@@
×$\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},s)$ $=s\varepsilon_{0}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (2)
@@ II-B Inhomogeneous Wave Equation for the Electric Field in the Laplace Transform Domain
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (5)
@@
where we used $\nicefrac{{1}}{{c^{2}}}=\mu_{0}\varepsilon_{0}$. Equation (5) is the inhomogenous wave equation for the electric field in the Laplace transform domain.
In order to obtain a more usable form, where we can incorporate common initial conditions, using Ampere’s law in the time domain, ×$\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t)=\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)+\mathbf{j}_{s}(\mathbf{r},t)\rightarrow\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t^{-})=\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)\left|{}_{t=t^{-}}\right.+\mathbf{j}_{s}(\mathbf{r},t^{-})$, and the electric constitutive relation in the time domain, $\mathbf{D}(\mathbf{r},t)=\varepsilon_{0}\mathbf{E}(\mathbf{r},t)+\mathbf{P}(\mathbf{r},t)\rightarrow\mathbf{P}(\mathbf{r},t^{-})=\mathbf{D}(\mathbf{r},t^{-})-\varepsilon_{0}\mathbf{E}(\mathbf{r},t^{-})$, we can write (5) as
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (6)
@@
We can use (6) to solve for the electric field in any temporal region, where the initial conditions are known at the temporal boundary immediately preceding the temporal region. We assume that there is no change in the induced current density just after the time-change, i.e., at $t=t_{\alpha}^{+}$, where $\alpha=0,1,2,...$. That is, although the material permittivity is modeled as changing value instantaneously, the material response (electrons) cannot change instantaneously due to their mass. This leads to the temporal boundary conditions (BCs); all components of the electric field $\mathbf{E}$, the magnetic field $\mathbf{H}$, and the current density $\mathbf{J}$ are continuous at the temporal boundary. Then the fields are the same just before and just after the time-change with respect to the media (and consequently at the time-change $t=t_{\alpha}$), i.e., at $t=t_{\alpha}^{-}$ and $t=t_{\alpha}^{+}$. In other words, we can use $t=t_{\alpha}^{-}$, $t=t_{\alpha}$, or $t=t_{\alpha}^{+}$ for the initial condition parameters.
@@ II-C One Temporal Boundary
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}(\mathbf{r},s)$ (7)
@@
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}(\mathbf{r},s)$ (8)
@@
which is the inhomogenous wave equation for the electric field in the Laplace transform domain (when used for spatial region 1 we set $\mathbf{j}_{s}(\mathbf{r})$ to zero).
We can use the inhomogenous wave equation for the electric field in the spatial regions in temporal region 1,
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r})$ $=\frac{\omega_{0}^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm1}(\mathbf{r},\omega_{0})\cdot\mathbf{E}_{m1}(\mathbf{r})$ (9)
@@ II-D Obtaining the Standard Form for the Wave Equation
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)-s\mu_{0}\mathbf{j}_{s}^{\prime}(\mathbf{r},s),$ (14)
@@ II-E Green’s Functions in the Laplace Transform Domain
(××$\left(\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times+\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\right)\underline{\mathbf{G}}_{m2}^{\prime}(\mathbf{r},\mathbf{r}^{\prime},s)$ $=\underline{\mathbf{I}}\delta(\mathbf{r}-\mathbf{r}^{\prime}),$ (22)
@@ II-E2 Isotropic Media
[…]
@@
[…]
@@
[…]
@@ II-E4 Fields
[…]
@@
where $C_{11}^{e}=i\omega_{0}\mu_{0}\left(\nicefrac{{\varepsilon_{11}}}{{\varepsilon_{21}}}\right)$, $C_{21}^{e}=i\omega_{0}\mu_{0}$, $C_{12}^{e}=A_{2}^{s}(s)\left(-s\mu_{0}\right)\left(\nicefrac{{\varepsilon_{12}(s)}}{{\varepsilon_{22}(s)}}\right)$, $C_{22}^{e}=A_{2}^{s}(s)\left(-s\mu_{0}\right)$, $C_{11}^{m}=\nicefrac{{\varepsilon_{11}}}{{\varepsilon_{21}}}$, $C_{21}^{m}=1$, $C_{12}^{m}=A_{2}^{s}(s)\left(\nicefrac{{\varepsilon_{12}(s)}}{{\varepsilon_{22}(s)}}\right)$, and $C_{22}^{m}=A_{2}^{s}(s)$
@@ III Results
In the following we obtain results for material configurations that give a strong SPP response. In all cases the results are obtained for the configuration in Fig. 1, where, as before, we assume nonmagnetic materials, piece-wise homogeneous layers, and isotropic media.
@@
$\ \varepsilon_{r21}(\omega_{0})=\varepsilon_{r21}^{\prime}+i\frac{\sigma_{21}}{\omega_{0}\varepsilon_{0}}\\ \\ \end{array}$ $\ \varepsilon_{r22}(s)=\varepsilon_{r22}^{\prime}+\frac{\sigma_{22}}{s\varepsilon_{0}}\\ \\ \end{array}$
$\ \varepsilon_{r11}(\omega_{0})=1-\frac{\omega_{p11}^{2}}{\omega_{0}^{2}+i\omega_{0}\Gamma_{d11}}\\ \\ \end{array}$ $\ \varepsilon_{r12}(s)=1+\frac{\omega_{p12}^{2}}{s^{2}+s\Gamma_{d12}}\\ \\ \end{array}$

### v2

@@ I Introduction
Lately, there has been a resurgence in this research topic [7, 6, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 10, 24, 25, 26, 27, 11, 28, 29, 12, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], driven by the potential for the applications in plasmonics (nano-photonics, nano-optics, photonic metamaterials), such as magnet-free nonreciprocity [15, 29], temporal aiming [21, 40], and extreme energy transformations [28], to name a few. There has been some work on time-varying media systems that support SPPs [16, 18, 22]; however, these deal with existing propagating SPPs encountering a temporal boundary. To the best of our knowledge, there has not been work on dipole excitation of SPPs in a time-varying system. This is the topic of this work [41], where such a system allows for the analysis of the interaction among dipole excitations, where we have the ability to consider the interactions at the moment of SPP creation (e.g., as a time-varying system changes from one that does not support SPPs to one that supports SPP propagation).
@@
In the subsequent sections we establish the theoretical framework for the specific configuration of a time-varying media system that supports SPPs, in at least one temporal region, or both, excited by a dipole source, where the media configuration suddenly changes at a temporal boundary. We first state Maxwell’s equations in the Laplace transform domain, derive the inhomogeneous electric field wave equation in the Laplace transform domain for generic media, i.e., linear, inhomogeneous, anisotropic, temporally dispersive (temporally nonlocal), and spatially nondispersive (local) media, and then illustrate the use of Green’s functions to solve the inhomogeneous wave equation to obtain the electric field.
@@
The model for this configuration, seen in Fig. 1, defines two temporal regions (one temporal boundary). This framework can be extended to the multi-temporal boundary case by solving the inhomogeneous electric field wave equation in the Laplace transform domain in each temporal region, using the Green’s functions applicable to each temporal region, where the initial conditions at the preceding temporal boundary are applicable.
Figure 1: Model for the general case of a single interface between two different materials ($\mu_{2}$$\varepsilon_{2}$ for $z>0$, $\mu_{1}$$\varepsilon_{1}$ for $z<0$) with an electric dipole point source, $\widehat{\mathbf{z}}j_{zs}$, located along the interface surface, where we consider two temporal regions (one temporal boundary). The double subscript notation is described in the text. Additionally, $q_{x}$, $q_{y}$, and $\phi$ define the coordinate system in the momentum space.
@@ II-A Maxwell’s Equations in the Laplace Transform Domain
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-s\mu_{0}\underline{\boldsymbol{\mu}}_{r}(\mathbf{r},s)\cdot\mathbf{H}(\mathbf{r},s)$ (1)
@@
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},s)$ $=s\varepsilon_{0}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (2)
@@ II-B Inhomogeneous Wave Equation for the Electric Field in the Laplace Transform Domain
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (5)
@@
where we used $\nicefrac{{1}}{{c^{2}}}=\mu_{0}\varepsilon_{0}$. Equation (5) is the inhomogeneous wave equation for the electric field in the Laplace transform domain.
In order to obtain a more usable form, where we can incorporate common initial conditions, using Ampere’s law in the time domain, $\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t)=\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)+\mathbf{j}_{s}(\mathbf{r},t)\rightarrow\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t^{-})=\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)\left|{}_{t=t^{-}}\right.+\mathbf{j}_{s}(\mathbf{r},t^{-})$, and the electric constitutive relation in the time domain, $\mathbf{D}(\mathbf{r},t)=\varepsilon_{0}\mathbf{E}(\mathbf{r},t)+\mathbf{P}(\mathbf{r},t)\rightarrow\mathbf{P}(\mathbf{r},t^{-})=\mathbf{D}(\mathbf{r},t^{-})-\varepsilon_{0}\mathbf{E}(\mathbf{r},t^{-})$, we can write (5) as
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (6)
@@
We can use (6) to solve for the electric field in any temporal region, where the initial conditions are known at the temporal boundary immediately preceding the temporal region. We assume that there is no change in the induced current density just after the time-change, i.e., at $t=t_{\alpha}^{+}$, where $\alpha=0,1,2,$.... That is, although the material permittivity is modeled as changing value instantaneously, the material response (electrons) cannot change instantaneously due to their mass. This leads to the temporal boundary conditions (BCs); all components of the electric field $\mathbf{E}$, the magnetic field $\mathbf{H}$, and the current density $\mathbf{J}$ are continuous at the temporal boundary. Then the fields are the same just before and just after the time-change with respect to the media (and consequently at the time-change $t=t_{\alpha}$), i.e., at $t=t_{\alpha}^{-}$ and $t=t_{\alpha}^{+}$. In other words, we can use $t_{\alpha}^{-}, t=t$$t_{\alpha}$, or $t=t_{\alpha}^{+}$ for the initial condition parameters.
@@ II-C One Temporal Boundary
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}(\mathbf{r},s)$ (7)
@@
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}(\mathbf{r},s)$ (8)
@@
which is the inhomogeneous wave equation for the electric field in the Laplace transform domain (when used for spatial region 1 we set $\mathbf{j}_{s}(\mathbf{r})$ to zero).
We can use the inhomogeneous wave equation for the electric field in the spatial regions in temporal region 1,
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r})$ $=\frac{\omega_{0}^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm1}(\mathbf{r},\omega_{0})\cdot\mathbf{E}_{m1}(\mathbf{r})$ (9)
@@ II-D Obtaining the Standard Form for the Wave Equation
$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)-s\mu_{0}\mathbf{j}_{s}^{\prime}(\mathbf{r},s),$ (14)
@@ II-E Green’s Functions in the Laplace Transform Domain
$\left(\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times+\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\right)\underline{\mathbf{G}}_{m2}^{\prime}(\mathbf{r},\mathbf{r}^{\prime},s)$ $=\underline{\mathbf{I}}\delta(\mathbf{r}-\mathbf{r}^{\prime}),$ (22)
@@ II-E2 Isotropic Media
$g_{zz2n}^{p}(\mathbf{r},\mathbf{r}^{\prime})$ $=\frac{1}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}dq_{x}dq_{y}$ (27)
@@
$g_{zz2n}^{s}(\mathbf{r},\mathbf{r}^{\prime})$ $=\frac{1}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}dq_{x}dq_{y}$ (28)
@@
$g_{zz1n}(\mathbf{r},\mathbf{r}^{\prime})$ $=\frac{1}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}dq_{x}dq_{y}$ (29)
@@ II-E4 Fields
For a vertical unit point dipole current source, $\mathbf{j}_{s}(\mathbf{r})=\widehat{\mathbf{z}}\delta(\mathbf{r}-\mathbf{r}_{0})$, using $\mathbf{E}_{m1}(\mathbf{r}),\>\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)=\left(k_{mn}^{2}+\mathbf{\nabla}\mathbf{\nabla}\cdot\right)\intop_{V}dV^{\prime}\underline{\mathbf{g}}_{mn}(\mathbf{r},\mathbf{r}^{\prime})\cdot\mathbf{S}_{n}(\mathbf{r}^{\prime})$ and $\mathbf{H}_{m1}(\mathbf{r}),\>\mathbf{H}_{m2}^{\prime}(\mathbf{r},s)=K_{mn}\mathbf{\nabla}\times\intop_{V}dV^{\prime}\underline{\mathbf{g}}_{mn}(\mathbf{r},\mathbf{r}^{\prime})\cdot\mathbf{S}_{n}(\mathbf{r}^{\prime})$, where $\mathbf{S}_{1}(\mathbf{r}^{\prime})=\nicefrac{{\mathbf{j}_{s}(\mathbf{r}^{\prime})}}{{\left(-i\omega_{0}\varepsilon_{21}\right)}}$, $\mathbf{S}_{2}(\mathbf{r}^{\prime})=\nicefrac{{\mathbf{j}_{s}^{\prime}(\mathbf{r}^{\prime},s)}}{{\left(s\varepsilon_{22}(s)\right)}}$, $K_{m1}=-i\omega_{0}\varepsilon_{m1}$, $K_{m2}=s\varepsilon_{m2}(s)$, and $\mathbf{j}_{s}^{\prime}(\mathbf{r}^{\prime},s)=A_{2}^{s}(s)\mathbf{j}_{s}(\mathbf{r}^{\prime})$, we can write the fields for each respective temporal region as
@@
where $C_{11}^{e}=i\omega_{0}\mu_{0}\left(\nicefrac{{\varepsilon_{11}}}{{\varepsilon_{21}}}\right)$, $C_{21}^{e}=i\omega_{0}\mu_{0}$, $C_{12}^{e}=A_{2}^{s}(s)\left(-s\mu_{0}\right)\left(\nicefrac{{\varepsilon_{12}(s)}}{{\varepsilon_{22}(s)}}\right)$, $C_{22}^{e}=A_{2}^{s}(s)\left(-s\mu_{0}\right)$, $C_{11}^{m}=\nicefrac{{\varepsilon_{11}}}{{\varepsilon_{21}}}$, $C_{21}^{m}=1$, $C_{12}^{m}=A_{2}^{s}(s)\left(\nicefrac{{\varepsilon_{12}(s)}}{{\varepsilon_{22}(s)}}\right)$, and $C_{22}^{m}=A_{2}^{s}(s)$ .
@@ III Results
In the following we obtain results for material configurations that give a strong SPP response. As a check on the validity of the model presented here, we obtained results for the configurations used in [5], which concerns a dipole source in a spatially-homogeneous time-changing medium. We found good agreement with the results (Figs. 1-3) in [5], with some small differences attributed to the different numerical methods used. We also obtained the same shifted frequency values determined in [5]. These comparisons provided confidence that the model is correct, since here we only change the Green’s function to account for inhomogeneous media and the presence of SPP-supporting interfaces. We note that the long-time transient field response approaches the steady-state field after the temporal-boundary transients have died out, which must occur, and leads to confidence in the method. Additionally, we observed the effects of the interference between the forward and backward waves (from the temporal-boundary transient SPPs), which aligns with the observations described in [16]. In that work, a propagating SPP is the wave incident on a temporal boundary (not the field of a dipole source as we consider here). We note that determining the shifted frequency values in this work becomes much more difficult since there are numerous modal temporal frequencies for the supported SPP modes. Further analysis regarding this, as well as other areas, e.g., energy conversion at the temporal boundary (potentially utilizing FDTD simulations), would be beneficial and a worthwhile pursuit for future work. Here, we focused on first establishing the necessary framework for further exploration of these areas. In all cases the results are obtained for the configuration in Fig. 1, where, as before, we assume nonmagnetic materials, piece-wise homogeneous layers, and isotropic media.
@@
[…]
@@
Note that in practice these configurations may be created for experimentation (practical application) by laser-induced plasma creation for the case of creating a sudden dense plasma channel in air [47, 48, 49, 50] (applicable to the configuration described in Table I), and, for the case of time-varying dielectric-plasma interfaces in both temporal regions (the configuration described in Table II), the method in [51] can be applied, where the material parameters of a SPP-supporting platform (at optical frequencies) are modulated in time by a fast-switching magnetic bias.
@@
[…]
l0ptb79pt(a)Refer to caption
[…]
l0ptb79pt(b)Refer to caption
[…]
l0ptb79pt(c)Refer to caption
[…]
l0ptb79pt(d)Refer to caption
[…]
l0ptb96pt(e)Refer to caption
[…]
l0ptb96pt(f)Refer to caption